two similar trapezoids have area 96 cm2 and 24 cm2. Find the ratio of similarity.
Answer by venugopalramana(3286)

AREA=0.5*(SUM OF PARALLEL SIDES)*HEIGHT=0.5(B+T)*H
IF RATIO OF PROPORTIONALITY =X WE HAVE
A1=0.5(B+T)*H.FOR TRAPEZIUM 1
A2=0.5(BX+TX)*HX=0.5(B+T)*H*X^2=A1*X^2
BUT A2/A1=96/24 GIVEN
X^2=96/24=4
X=2